OpenStudy (anonymous):

if sin 53 degrees is close to 4/5, which is closest to the length of NO?

OpenStudy (anonymous):

theres a right triangle at the top that got cut off

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

a. 4 b.80 c. 20 d.10

hartnn (hartnn):

sin = opposite side / hypotenuse

hartnn (hartnn):

so, sin 53 = NO/ 100

hartnn (hartnn):

4/5 = NO / 100 find NO from here

OpenStudy (anonymous):

would the answer be 80?

OpenStudy (anonymous):

because 4/5 of 100 is 80

hartnn (hartnn):

80 is correct :)

OpenStudy (anonymous):

Yay!!!! thank you thank you thank you

hartnn (hartnn):

welcome ^_^ happy to help :)

OpenStudy (anonymous):

okay heres my last one!

hartnn (hartnn):

a = amplitude = 4 period = 2pi/ b = pi/2 find b from here

OpenStudy (anonymous):

would b be 4?

hartnn (hartnn):

correct! so you have the equation now

OpenStudy (anonymous):

is the answer b ? i'm not sure

OpenStudy (anonymous):

i take it back it's C!

hartnn (hartnn):

y = - 4 sin 4 theta

hartnn (hartnn):

yes! correct :)